# Diffusion Models in Simulation-Based Inference: A Tutorial Review

### v1

@@ 6.4 Case Study 4: Pooling Regimes in Cognitive Modeling
In our final case study, we illustrate the utility of compositional inference for different structured Bayesian models (Section 4.2). Our simulator is an evidence accumulation model (EAM) [ratcliff2016diffusion], widely used to study decision-making and recently used to benchmark amortized hierarchical inference [habermann2024amortized]. The EAM describes binary decision-making as the accumulation of noisy information until one of two boundaries is reached (Figure 13A). Each trial is defined by four parameters: drift rate $\nu$, decision threshold $\alpha$, non-decision time $t_{0}$, and starting bias $\beta$. We simulate latent trajectories using the Euler–Maruyama scheme with absorbing boundaries at 0 and $\alpha$. The simulation returns both a binary choice and a reaction time (RT), with trials censored at $t=\text{max}_{t}$ if no boundary is hit (Section A.1.4).
For this case study, we have two different priors: a flat and a hierarchical one. For each, we train a diffusion model with EDM noise schedule. For training, we simulate training instances with $R=1$ subject and $N=30$ trials per subject. Hence, as a summary network, we employ a SetTransformer [lee2019set] to account for the exchangeability of the trials. For inference, we simulate 100 test sets of $R=100$ subjects, each with $N=30$ trials. Thus, for the same subject, we have repeated observations of trials generated by the same parameter and, depending on the prior, parameters between subjects are assumed to be the same (complete pooling) or linked by a hierarchical structure (partial pooling). Details of the simulator and priors are given in Section A.1.4.
@@
Here $M$ is the size of the mini-batch and $d(t)$ a time-dependent damping factor to reduce the error aggregation of the compositional score, which is necessary as naive summation of the subject scores becomes unstable already for more than 30 subjects [arruda2025compositional]. This model yields a sharper global posterior by training only on single-subject simulations. As shown by the NRMSE in Figure 13B, additional evidence across subjects substantially improves estimation. In contrast, direct training without composition would require $R$ times as many simulations per training instance and would not scale to large data sets [arruda2025compositional].
@@ 7 Discussion
Our empirical evaluations across the first three case studies reveal that design choices matter substantially, but their impact is dimension-dependent. For the low-dimensional benchmark problems (Case Study 1, see Section 6.1), diffusion models with a variance-preserving EDM schedule were the best choice. For high-dimensional inference (Case Study 2, see Section 6.2), flow matching—particularly with optimal transport coupling—achieved the best overall performance in terms of normalized RMSE and calibration error. The variance-preserving EDM schedule with $\mathbf{F}$-prediction provided a competitive alternative, while the variance-exploding EDM schedule exhibited higher calibration errors in this setting.
These results align with prior observations that variance-preserving schedules generally perform better than VE schedules for high-dimensional targets [SongSoh2020]. We observe analogous trends in the GRF experiment (Case Study 3, see Section 6.3). In the low-dimensional setting, variance-preserving EDM-based diffusion models usually perform best, followed closely by flow-matching and consistency models. In the high-dimensional setting, flow matching is markedly more sensitive to the available simulation budget, whereas the diffusion model remains competitive even under tight budgets. By contrast, the consistency model performs poorly across simulation budgets and scales in this setting. This presumably relates to the additional burden of approximating the denoising trajectory in consistency models under a fixed backbone, which appears particularly challenging when the output is high-dimensional.
@@
In practice, we recommend using diffusion models with the variance-preserving EDM noise schedule and an adaptive SDE solver or flow matching with an adaptive higher-order ODE solver, as these appear to be the best-performing models across SBI problems. For large problems, the advantage of consistency models in terms of inference speed becomes more pronounced and therefore may provide a good starting point for practitioners. Promising new directions include energy matching [balcerak2025energy] and equilibrium matching [wang2025equilibrium] models, which learn a time-independent flow field, thereby simplifying the architecture of the backbone and outperforming flow-based models in image generation quality.
@@
Our analysis of pooling regimes (Case Study 4, see Section 6.4) demonstrates the substantial efficiency gains elicited by compositional score matching. For the target cognitive model, compositional estimation reduced the simulation budget by multiple orders of magnitude compared to naive training. Furthermore, compositional estimation is applicable to hierarchical Bayesian models, reducing uncertainty in local parameters. However, we also observed that naive compositional estimation can suffer from error accumulation, necessitating some form of error-damping or stabilization [arruda2025compositional].
@@
The fundamental question remains: under what conditions does guidance preserve statistical accuracy? Recent work by [vuong2025we] suggests that samples will align with the marginals implied by the guidance, even if intermediate trajectories differ from true reverse diffusion. However, extensive characterization of the conditions under which this holds represents a critical open problem. This is particularly pressing for compositional estimation [geffner2023compositional, linhart2024diffusion, gloeckler2025compositional, arruda2025compositional], where score composition across many observations amplifies guidance-induced errors.
@@
Taken together, these findings suggest that C2ST can be useful but fragile, and it also comes with additional design choices: one must decide how many and which samples to use for training the classifier. This is an expensive step in high-dimensional SBI, where sampling from multi-step generative models is costly, and the classifier architecture itself has to be adapted to the data structure, adding hyperparameters and complicating comparisons across studies. Developing calibration diagnostics that remain informative in high dimensions and can distinguish approximation error from misspecification represents an urgent methodological need. Recent methods such as posterior SBC [sailynoja2025posterior] and alternative test quantities for SBC [modrak2025simulation] offer promising directions, but their computational cost and sensitivity in extreme dimensions require further improvements.
Even though diffusion models enable amortized inference with fixed simulation budgets, many scientific domains face severe constraints where generating even thousands of high-fidelity simulations remains infeasible [Ohana2024-well]. In fields such as fluid dynamics, producing a high-fidelity pair $(\boldsymbol{\theta},\mathbf{y}_{\text{obs}})$ can take multiple days, rendering simulation-based training hardly feasible. Our review highlights a critical infrastructure gap: most SBI papers still evaluate exclusively on low-dimensional benchmarks, and applications to high-dimensional parameters remain rare. The field would benefit enormously from expanded benchmark suites with controllable dimensionality regimes and simulation budgets.
@@
Diffusion models offer unique flexibility through score composition and guidance, which may be leveraged to mitigate the effects of model misspecification. This is especially relevant in real-world applications, which invariably involve some form of misspecification, making the “Sim2Real” problem hard [schmitt2023detecting]. Recent work explores robust inference through generalized Bayesian approaches [gao2023generalized], unsupervised domain adaptation [elsemuller2025does], iterative refinement for misspecified priors [barco2025tackling], or leverage of scarce ground-truth measurements to correct for misspecification in flow matching [ruhlmann2025flow]. Yet, diffusion models remain largely unexplored in the context of misspecified models [kelly2025simulation]. Can we design objectives that explicitly trade off fidelity to the (potentially misspecified) simulator against empirical data fit? How should we combine scores from multiple competing models to achieve robust inference? These questions connect to broader themes in Bayesian model averaging and ensemble methods but require rethinking in the SBI context.
@@ Acknowledgments
This work was supported by the European Union via the ERC grant INTEGRATE, grant agreement number 101126146, and under Germany’s Excellence Strategy by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) (EXC 2047-390685813, EXC 2151-390873048, and 524747443), by the German Federal Ministry of Education and Research (BMBF) through the EMUNE project (031L0293A), by the Klaus Tschira Stiftung for their support via the SIMPLAIX project, and the University of Bonn via the Schlegel Professorship of J.H.. STR and NB are supported by the National Science Foundation under Grant No. 2448380. We thank Jerry Huang and Lasse Elsemüller for their support in creating figures 3 and 13.
@@ A.1 Empirical Evaluation: Additional Information
All models and samplers were implemented in BayesFlow 2.0.7 [bayesflow_2023_software]. We use z-score normalization as a target normalization strategy, as it was shown to give superior performance compared to min-max scaling [orsini2025flow2]. In our EDM settings, we adopt $\sigma_{\text{target}}=1$, as we standardize our target during training, and we set $\sigma_{\text{min}}=0.0001$ and use the EDM weighting. For the cosine schedule, we use no shift $s=0$ and a VP schedule with sigmoid weighting. For the consistency models, we employ multi-step sampling during inference with 10 (discrete case) or 15 steps (continuous case). All models were trained using Adam (for online training) or AdamW (for offline training) with a cosine schedule for the learning rate with an initial value of 5e-4.
A compact multilayer perceptron serves as the core backbone for all models. It consists of five fully connected layers with width 256, Mish activations, He normal initialization, and residual connections. A small dropout rate of 0.05 is used for flow matching and consistency models and no dropout for the diffusion models.
@@
where $\beta_{i}=(1-1/i)^{\gamma+1}$ and we set $\gamma=6.94$ as discussed in [karras2024analyzing]. The averaged weights are then used during inference, which often leads to better generalization during inference with diffusion models [karras2024analyzing].
@@
Following [lopez-paz2017revisiting], we quantify the discrepancy between an approximate posterior $q(\boldsymbol{\theta}\mid\mathbf{y})$ and a reference (“ground-truth”) posterior $p(\boldsymbol{\theta}\mid\mathbf{y})$ via a classifier two-sample test. For a given observation $\mathbf{y}$, we draw i.i.d. samples
@@
Without access to reference posterior samples, when can leverage C2ST also using samples from the prior alongside with simulations and posterior samples for the corresponding simulation [yao2023discriminative], which we used in our third case study. In our second case study, the classifier takes as input both parameters and the corresponding log-likelihood values as sufficient summary statistic of the data.
@@
Maximum mean discrepancy provides a kernel-based, nonparametric measure of discrepancy between two distributions [gretton2012kernel]. Given posterior samples from an approximate distribution $\hat{\boldsymbol{\theta}}\sim q(\boldsymbol{\theta}\mid\mathbf{y}_{r})$ and reference samples from a reference (“ground-truth”) posterior $\boldsymbol{\theta}^{\star}\sim p(\boldsymbol{\theta}\mid\mathbf{y}_{r})$, the squared MMD is defined as
@@ A.1.1 Case Study 1: Low-Dimensional Benchmarks
We ran all benchmark problems using online training with 1,000 epochs, a batch size of 128 and 32,000 simulations. We then sampled as many posterior samples as in the reference posterior for each observation in the benchmark collection and compared the model with the corresponding sampler using C2ST configured as provided by [lueckmann2021benchmarking]. For each model, we tested different ODE and SDE (if possible) and timed the inference time. For each sampler, we used 500 steps or used an adaptive method with minimal 50 steps and maximal 1,000 steps (which was never reached). As classifier for the C2ST, we use the setup provided by [lueckmann2021benchmarking].
@@
For a detailed description of the prior and simulators, we refer to the original benchmark [lueckmann2021benchmarking]. To improve numerical stability and enforce parameter constraints, we apply simple parameter transformations depending on the problem. For the Lotka-Volterra and SIR models, a log transformation is applied to the parameters to enforce positivity. For the Gaussian mixture ([10,10][-10,10]), SLCP with and without distractors ([3,3][-3,3]), Gaussian linear with uniform prior ([1,1][-1,1]), and two-moons problems ([1,1][-1,1]), parameters are constrained to the corresponding fixed bounded interval via a sigmoid function. All parameters are transformed for training and back-transformed after posterior sampling. Transformed parameters and observables are standardized during training.
@@
The ODE samplers included Euler, RK45 (with the coefficients used in the Dormand–Prince variant [dormand1996numerical], which is the standard ODE solver in many libraries), and TSIT5 (which has newer coefficients for Runge–Kutta that seem to be more efficient [tsitouras2011runge]), in both fixed-step and adaptive variants. For the adaptive variants, we used the embedded 4th-order solution to form a scalar error estimate by taking the maximum $\ell_{2}$-norm over all state components, and updated the step size via a standard controller $h_{\text{new}}=h\cdot 0.9(\text{tol}/(\text{err}+10^{-12}))^{1/5}$, clipped to stay within a factor of $[0.2,5]$ and constrained so that the total steps remained between 50 and 1,000 (default tolerance $10^{-6}$).
@@
The SDE samplers covered Euler–Maruyama (EulerM) (with fixed, adaptive, and predictor–corrector variants), SEA (an improved one-step method with strong order 1 for SDEs with additive noise) [foster2024high], ShARK [foster2024high] (two-step method with strong order 1.5 for SDEs with additive noise), an adaptive two-step method [Jolicoeur-MartineauLi2021], and annealed Langevin dynamics [SongErm2020a, song2020improved]. The Euler–Maruyama adaptive variant followed [fang2020adaptive], setting the step size proportional to $\max(1,\|\mathbf{x}\|^{2})/\max(1,\|\mathbf{v}(\mathbf{x})\|^{2})$ with clipping based on minimal/maximal steps. The two-step adaptive scheme of [Jolicoeur-MartineauLi2021] was implemented as an Euler–Heun predictor–corrector: we estimated the local error from the discrepancy between the Euler and Heun updates, compared the normalized error against combined relative/absolute tolerances ($\epsilon_{\text{abs}}=0.02576$, $1\%$ of $99\%$ CI of standard normal as targets are standardized, and $\epsilon_{\text{rel}}=0.1$), and used an accept–reject step with settings from [Jolicoeur-MartineauLi2021] for step-size adaptation. Step size was also clipped based on minimal/maximal allowed steps. The Langevin-based sampler used four times the nominal step count, and predictor–corrector (PC) versions applied one (Euler–Maruyama) or five (Langevin) corrector updates per step. The implementation of the Langevin sampler was based on recommendations in [song2020improved] and for PC based on [SongSoh2020].
@@
Optimal transport flow matching replaces the independent pairing of samples from the base and target distributions with a coupling induced by an (entropically regularized) optimal transport plan [tong2024improving, pooladian2023multisample]. Given a mini-batch $\{\mathbf{z}_{0}^{(i)},\mathbf{z}_{1}^{(i)}\}_{i=1}^{n}$, a cost matrix $\mathbf{C}_{ij}=\|\mathbf{z}_{0}^{(i)}-\mathbf{z}_{1}^{(j)}\|_{2}^{2}$ is constructed and converted into a kernel
@@
where $\epsilon>0$ is an entropic regularization parameter. The Sinkhorn–Knopp algorithm then alternates row and column normalizations of $\mathbf{K}$ until the prescribed marginals are matched, yielding a transport plan $\pi\in\mathbb{R}^{n\times m}$ [sinkhorn1967concerning, cuturi2013sinkhorn]. We set $\epsilon=0.1$, a maximal number of 100 optimization steps and use a log-stabilized version of the Sinkhorn–Knopp algorithm.
While balanced uniform marginals are used for standard OT, partial optimal transport augments the problem with a dummy mass to transport only a fraction $s\in(0,1)$ of the probability mass, which can improve robustness to misspecified pairings in mini-batch settings [nguyen2022improving]. For conditional OT, the transport cost can be modified by incorporating a condition-dependent penalty, effectively restricting transport to pairs with similar conditions and reducing spurious matches [Fluri2024-improvingFlowMatchingSimulation-basedInferenceu, cheng2025curse]. We compute a pairwise cosine distance for all conditions and add this to the cost matrix $\mathbf{C}$ with a weight $w$ determined by the ratio $r(w)$, which measures the proportion of samples that are considered potential optimal transport candidates as described by [cheng2025curse].
We benchmarked standard optimal transport (OT), partial optimal transport (POT) with $s=0.8$, conditional optimal transport (COT) with a $r(w)=0.01$ and both strategies together (CPOT) with $r(w)=0.01$ and $s=0.9$. As flow matching with standard optimal transport performed rather poorly on the low-dimensional tasks, we additionally rerun the problems with a fixed simulation budget of 32,000 simulations. The performance in terms of C2ST was comparable to online training. We generally observe that not enough contraction is achieved with standard or partial OT. COT, and CPOT all improved upon standard OT, reaching a similar level of accuracy as without any form of OT (see Figure A.2).
Figure A.1: Extended Case Study. Mean (and std) of C2ST and timing for all samplers across all benchmark problems. For C2ST we excluded here the Lotka-Volterra task to reduce the overall variance.
Figure A.2: Optimal Transport Flow Matching. Mean (and std) of C2ST for all variants of optimal transport flow matching across all benchmark problems.

### v2

@@ 6.4 Case Study 4: Pooling Regimes in Cognitive Modeling
In our final case study, we illustrate the utility of compositional inference for different structured Bayesian models (Section 4.2). Our simulator is an evidence accumulation model (EAM) [141], widely used to study decision-making and recently used to benchmark amortized hierarchical inference [65]. The EAM describes binary decision-making as the accumulation of noisy information until one of two boundaries is reached (Figure 13A). Each trial is defined by four parameters: drift rate $\nu$, decision threshold $\alpha$, non-decision time $t_{0}$, and starting bias $\beta$. We simulate latent trajectories using the Euler–Maruyama scheme with absorbing boundaries at 0 and $\alpha$. The simulation returns both a binary choice and a reaction time (RT), with trials censored at $t=\text{max}_{t}$ if no boundary is hit (Section A.1.4).
For this case study, we have two different priors: a flat and a hierarchical one. For each, we train a diffusion model with EDM noise schedule. For training, we simulate training instances with $R=1$ subject and $N=30$ trials per subject. Hence, as a summary network, we employ a SetTransformer [99] to account for the exchangeability of the trials. For inference, we simulate 100 test sets of $R=100$ subjects, each with $N=30$ trials. Thus, for the same subject, we have repeated observations of trials generated by the same parameter and, depending on the prior, parameters between subjects are assumed to be the same (complete pooling) or linked by a hierarchical structure (partial pooling). Details of the simulator and priors are given in Section A.1.4.
@@
Here $M$ is the size of the mini-batch and $d(t)$ a time-dependent damping factor to reduce the error aggregation of the compositional score, which is necessary as naive summation of the subject scores becomes unstable already for more than 30 subjects [4]. This model yields a sharper global posterior by training only on single-subject simulations. As shown by the NRMSE in Figure 13B, additional evidence across subjects substantially improves estimation. In contrast, direct training without composition would require $R$ times as many simulations per training instance and would not scale to large data sets [4].
@@ 7 Discussion
Our empirical evaluations across the first three case studies reveal that design choices matter substantially, but their impact is dimension-dependent. For the low-dimensional benchmark problems (Case Study 1, see Section 6.1), diffusion models with a variance-preserving EDM schedule were the best choice. For high-dimensional inference (Case Study 2, see Section 6.2), flow matching—particularly with optimal transport coupling—achieved the best overall performance in terms of normalized RMSE and calibration error. The variance-preserving cosine schedule with $\mathbf{v}$-prediction provided a competitive alternative, while the variance-exploding EDM schedule exhibited higher calibration errors in this setting.
These results align with prior observations that variance-preserving schedules generally perform better than VE schedules for high-dimensional targets [166]. We observe analogous trends in the GRF experiment (Case Study 3, see Section 6.3). In the low-dimensional setting, variance-preserving EDM-based diffusion models usually perform best, followed closely by flow-matching and consistency models. In the high-dimensional setting, flow matching is markedly more sensitive to the available simulation budget, whereas the diffusion model remains competitive even under tight budgets. By contrast, the consistency model performs poorly across simulation budgets and scales in this setting. This presumably relates to the additional burden of approximating the denoising trajectory in consistency models under a fixed backbone, which appears particularly challenging when the output is high-dimensional.
@@
In practice, we recommend using diffusion models with the variance-preserving EDM noise schedule and an adaptive SDE solver or flow matching with an adaptive higher-order ODE solver, as these appear to be the best-performing models across SBI problems. For large problems, the advantage of consistency models in terms of inference speed becomes more pronounced and therefore may provide a good starting point for practitioners. Promising new directions include energy matching [7] and equilibrium matching [188] models, which learn a time-independent flow field, thereby simplifying the architecture of the backbone and outperforming flow-based models in image generation quality.
@@
Our analysis of pooling regimes (Case Study 4, see Section 6.4) demonstrates the substantial efficiency gains elicited by compositional score matching. For the target cognitive model, compositional estimation reduced the simulation budget by multiple orders of magnitude compared to naive training. Furthermore, compositional estimation is applicable to hierarchical Bayesian models, reducing uncertainty in local parameters. However, we also observed that naive compositional estimation can suffer from error accumulation, necessitating some form of error-damping or stabilization [4].
@@
The fundamental question remains: under what conditions does guidance preserve statistical accuracy? Recent work by [184] suggests that samples will align with the marginals implied by the guidance, even if intermediate trajectories differ from true reverse diffusion. However, extensive characterization of the conditions under which this holds represents a critical open problem. This is particularly pressing for compositional estimation [53, 105, 59, 4], where score composition across many observations amplifies guidance-induced errors.
@@
Taken together, these findings suggest that C2ST can be useful but fragile, and it also comes with additional design choices: one must decide how many and which samples to use for training the classifier. This is an expensive step in high-dimensional SBI, where sampling from multi-step generative models is costly, and the classifier architecture itself has to be adapted to the data structure, adding hyperparameters and complicating comparisons across studies. Developing calibration diagnostics that remain informative in high dimensions and can distinguish approximation error from misspecification represents an urgent methodological need. Recent methods such as posterior SBC [146] and alternative test quantities for SBC [118] offer promising directions, but their computational cost and sensitivity in extreme dimensions require further improvements.
Even though diffusion models enable amortized inference with fixed simulation budgets, many scientific domains face severe constraints where generating even thousands of high-fidelity simulations remains infeasible [124]. In fields such as fluid dynamics, producing a high-fidelity pair $(\boldsymbol{\theta},\mathbf{y}_{\text{obs}})$ can take multiple days, rendering simulation-based training hardly feasible. Our review highlights a critical infrastructure gap: most SBI papers still evaluate exclusively on low-dimensional benchmarks, and applications to high-dimensional parameters remain rare. The field would benefit enormously from expanded benchmark suites with controllable dimensionality regimes and simulation budgets.
@@
Diffusion models offer unique flexibility through score composition and guidance, which may be leveraged to mitigate the effects of model misspecification. This is especially relevant in real-world applications, which invariably involve some form of misspecification, making the “Sim2Real” problem hard [150]. Recent work explores robust inference through generalized Bayesian approaches [51], unsupervised domain adaptation [43], iterative refinement for misspecified priors [11], or leverage of scarce ground-truth measurements to correct for misspecification in flow matching [144]. Yet, diffusion models remain largely unexplored in the context of misspecified models [90]. Can we design objectives that explicitly trade off fidelity to the (potentially misspecified) simulator against empirical data fit? How should we combine scores from multiple competing models to achieve robust inference? These questions connect to broader themes in Bayesian model averaging and ensemble methods but require rethinking in the SBI context.
@@ Acknowledgments
This work was supported by the European Union via the ERC grant INTEGRATE, grant agreement number 101126146, and under Germany’s Excellence Strategy by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) (EXC 2047-390685813, EXC 2151-390873048, and 524747443), by the German Federal Ministry of Education and Research (BMBF) through the EMUNE project (031L0293A), by the Klaus Tschira Stiftung for their support via the SIMPLAIX project, and the University of Bonn via the Schlegel Professorship of JH. STR and NB are supported by the National Science Foundation under Grant No. 2448380. We thank Jerry Huang and Lasse Elsemüller for their support in creating figures 3 and 13.
@@ A.1 Empirical Evaluation: Additional Information
All models and samplers were implemented in BayesFlow 2.0.8 [140]. We use z-score normalization as a target normalization strategy, as it was shown to give superior performance compared to min-max scaling [127]. In our EDM settings, we adopt $\sigma_{\text{target}}=1$, as we standardize our target during training, and we set $\sigma_{\text{min}}=0.0001$ and use the EDM weighting. For the cosine schedule, we use no shift $s=0$ and a variance-preserving schedule with sigmoid weighting. We set $\lambda_{\mathrm{min}}=-15$ and $\lambda_{\mathrm{max}}=15$ and rescale the time accordingly. For the consistency models, we employ multi-step sampling during inference with 10 (discrete case) or 15 steps (continuous case). All models were trained using Adam (for online training) [94] or AdamW (for offline training) [112] with a cosine schedule for the learning rate [111] with an initial value of 5e-4.
A compact multilayer perceptron serves as the core backbone for all models. It consists of five fully connected layers with width 256, Mish activations [117], He normal initialization [68], and residual connections [69]. Targets and conditions are concatenated together and mapped to a shared embedding, time/SNR is embedded using a Fourier embedding [170] of dimension 32 similar to [190], while the residual blocks use a linear FiLM conditioning [134] for the embedded time and layer-wise normalization [6] after the skip-connection. A small dropout rate [167] of 0.05 is used for all models.
@@
where $\beta_{i}=(1-1/i)^{\gamma+1}$ and we set $\gamma=6.94$ as discussed in [89]. The averaged weights are then used during inference, which can lead to better generalization during inference with diffusion models [89].
@@
Following [110], we quantify the discrepancy between an approximate posterior $q(\boldsymbol{\theta}\mid\mathbf{y})$ and a reference (“ground-truth”) posterior $p(\boldsymbol{\theta}\mid\mathbf{y})$ via a classifier two-sample test. For a given observation $\mathbf{y}$, we draw i.i.d. samples
@@
Without access to reference posterior samples, when can leverage C2ST also using samples from the prior alongside with simulations and posterior samples for the corresponding simulation [193], which we used in our third case study. In our second case study, the classifier takes as input both parameters and the corresponding log-likelihood values as sufficient summary statistic of the data.
@@
Maximum mean discrepancy provides a kernel-based, nonparametric measure of discrepancy between two distributions [61]. Given posterior samples from an approximate distribution $\hat{\boldsymbol{\theta}}\sim q(\boldsymbol{\theta}\mid\mathbf{y}_{r})$ and reference samples from a reference (“ground-truth”) posterior $\boldsymbol{\theta}^{\star}\sim p(\boldsymbol{\theta}\mid\mathbf{y}_{r})$, the squared MMD is defined as
@@ A.1.1 Case Study 1: Low-Dimensional Benchmarks
We ran all benchmark problems using online training with 1,000 epochs, a batch size of 128 and 32,000 simulations. We then sampled as many posterior samples as in the reference posterior for each observation in the benchmark collection and compared the model with the corresponding sampler using C2ST configured as provided by [114]. For each model, we tested different ODE and SDE (if possible) and timed the inference time. For each sampler, we used 500 steps or used an adaptive method with minimal 50 steps and maximal 1,000 steps (which was never reached). As classifier for the C2ST, we use the setup provided by [114].
@@
For a detailed description of the prior and simulators, we refer to the original benchmark [114]. To improve numerical stability and enforce parameter constraints, we apply simple parameter transformations depending on the problem. For the Lotka-Volterra and SIR models, a log transformation is applied to the parameters to enforce positivity. For the Gaussian mixture ([10,10][-10,10]), SLCP with and without distractors ([3,3][-3,3]), Gaussian linear with uniform prior ([1,1][-1,1]), and two-moons problems ([1,1][-1,1]), parameters are constrained to the corresponding fixed bounded interval via a sigmoid function. All parameters are transformed for training and back-transformed after posterior sampling. Transformed parameters and observables are standardized during training.
@@
The ODE samplers included Euler, RK45 (with the coefficients used in the Dormand–Prince variant [41], which is the standard ODE solver in many libraries), and TSIT5 (which has newer coefficients for Runge–Kutta that seem to be more efficient [178]), in both fixed-step and adaptive variants. For the adaptive variants, we used the embedded 4th-order solution to form a scalar error estimate by taking the maximum $\ell_{2}$-norm over all state components, and updated the step size via a standard controller $h_{\text{new}}=h\cdot 0.9\bigl(\mathrm{err}/(\mathrm{atol}+\mathrm{rtol}\,\max_{i}|\mathbf{z}_{i}|)+10^{-12}\bigr)^{-1/5}$, clipped to stay within a factor of $[0.2,5]$ and further constrained by minimum and maximum step-size bounds (default $\mathrm{atol}=10^{-6}$, $\mathrm{rtol}=10^{-4}$).
@@
The SDE samplers covered Euler–Maruyama (EulerM) (with fixed, scheduled, adaptive, and predictor–corrector variants), SEA (an improved one-step method with strong order 1 for SDEs with additive noise) [48], ShARK [48] (two-step method with strong order 1.5 for SDEs with additive noise), an adaptive two-step method [87], and annealed Langevin dynamics [164, 165]. The Euler–Maruyama adaptive variant followed [45], setting the step size proportional to $\max(1,\|\mathbf{x}\|^{2})/\max(1,\|\mathbf{v}(\mathbf{x})\|^{2})$ with clipping based on minimal/maximal steps. The scheduled variant, uses a fixed step size in the log SNR. The two-step adaptive scheme of [87] was implemented as an Euler–Heun predictor–corrector: we estimated the local error from the discrepancy between the Euler and Heun updates, compared the normalized error against combined relative/absolute tolerances ($\epsilon_{\text{abs}}=0.02576$, $1\%$ of $99\%$ CI of standard normal as targets are standardized, and $\epsilon_{\text{rel}}=0.1$), and used an accept–reject step with settings from [87] for step-size adaptation. Step size was also clipped based on minimal/maximal allowed steps. The Langevin-based sampler used four times the nominal step count, and predictor–corrector (PC) versions applied one (Euler–Maruyama) or five (Langevin) corrector updates per step. The implementation of the Langevin sampler was based on recommendations in [165] and for PC based on [166].
@@
Optimal transport flow matching replaces the independent pairing of samples from the base and target distributions with a coupling induced by an (entropically regularized) optimal transport plan [175, 137]. Given a mini-batch $\{\mathbf{z}_{0}^{(i)},\mathbf{z}_{1}^{(i)}\}_{i=1}^{n}$, a cost matrix $\mathbf{C}_{ij}=\|\mathbf{z}_{0}^{(i)}-\mathbf{z}_{1}^{(j)}\|_{2}^{2}$ is constructed and converted into a kernel
@@
where $\epsilon>0$ is an entropic regularization parameter. The Sinkhorn–Knopp algorithm then alternates row and column normalizations of $\mathbf{K}$ until the prescribed marginals are matched, yielding a transport plan $\pi\in\mathbb{R}^{n\times m}$ [157, 32]. We set $\epsilon=0.1$, a maximal number of 100 optimization steps and use a log-stabilized version of the Sinkhorn–Knopp algorithm.
While balanced uniform marginals are used for standard OT, partial optimal transport augments the problem with a dummy mass to transport only a fraction $s\in(0,1)$ of the probability mass, which can improve robustness to misspecified pairings in mini-batch settings [122]. For conditional OT, the transport cost can be modified by incorporating a condition-dependent penalty, effectively restricting transport to pairs with similar conditions and reducing spurious matches [47, 26]. We compute a pairwise cosine distance for all conditions and add this to the cost matrix $\mathbf{C}$ with a weight $w$ determined by the ratio $r(w)$, which measures the proportion of samples that are considered potential optimal transport candidates as described by [26].
We benchmarked standard optimal transport (OT), partial optimal transport (POT) with $s=0.8$, conditional optimal transport (COT) with a $r(w)=0.01$ and both strategies together (CPOT) with $r(w)=0.01$ and $s=0.9$. We generally observe that not enough contraction is achieved with standard or partial OT. COT, and CPOT with $r(w)=0.01$ improved upon standard OT, reaching a similar level of accuracy as without any form of OT (see Figure A.2). Higher values of $r$ lead to straighter path, which can be beneficial if fewer steps in the sampler must be used, but straighter paths do not improve C2ST in general (see Figure A.2), e.g., for the SIR problem, straighter paths help while sampling with only 10 steps but not for more steps.
Figure A.1: Extended Case Study. Mean (and std) of C2ST and timing for all samplers across all benchmark problems.
Figure A.2: Optimal Transport Flow Matching. Mean (and std) of C2ST and timing for all variants of optimal transport flow matching across all benchmark problems.